OpenStudy (anonymous):

What is the vertex of the graph of y = –4(x + 2)2 + 5? (2, 5) (–2, 5) (5, –2) (5, 2)

OpenStudy (anonymous):

5, 2

OpenStudy (phi):

when a graph is in this special form (it is a parabola), the x value of the vertex is where (x-2) is zero. That is x= 2. And when (x-2) is zero, y= 0+5= 5, so the y value is 5 (2,5) is the vertex.